Calculate the client’s fluid maintenance? Calculate the IV rate for hourly maintenance fluids.

  1. Pediatric HOURLY fluid maintenance intake – based on weight in kg          

Calculate the fluid maintenance rate for your child who weigh 18 kg.?   

_____________________ ml/hour

                                  FORMULA:

a. Weight 1- 10 kg                                4 ml/kg for first 10 kg

b. Weight 10-20 kg                              40 mL + 2 ml/kg for next 10 kg (between 10 - 20 kg)

c. Weight >20 kg                                 60 + 1 ml/kg for remaining kg > 20 kg

  • Multiple total amount by 24 hours to determine 24 hour fluid requirement
Answer #1

Child weighs only 18 kg

Fluid maintenance rate of this 18 kg child=

According to formula, 40 ml + (2ml*8kg) =40 +16

=56 ml/hour

If, for 1 hour fluid maintenance is 56ml/hour then for 24 hour = 56*24 = 1344 ml/hour